OpenStudy (anonymous):

Simplify completely quantity x squared minus 100 over quantity x squared plus 2 x minus 120 divided quantity 6 x plus 60 over quantity x minus 2. I have done most of it, just need help with the end

sammixboo (sammixboo):

What do you have?

OpenStudy (anonymous):

let me type it out :)

sammixboo (sammixboo):

OK =) Tag me @sammixboo when you finish

OpenStudy (anonymous):

x^2 -100 x-2 ------------ * ------- = x^2 + 2x -120 6x + 60 (x - 10)(x+10) (x-2) ------------- * ------- = (x + 12)(x + 10) 6(x+10)

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

(x-10) (x-2) ------ * ________ (x+12) 6

OpenStudy (anonymous):

@sammixboo